On a math test with 52 questions , Kim solved 3/4 of the questions correctly. How many questions did Kim solve incorrectly ?
Answer by checkley79(3341)

3/4*52=3*52/4=156/4=39 CORRECTLY ANSWERED QUESTIONS.
52-39=13 INCORRECTLY ANSWERED QUESTIONS.